OpenStudy (anonymous):

what does mean that the function 1+3sin^4(2x) is going to be bounded since the sine is never larger than 1 ?

OpenStudy (anonymous):

does it mean that at the roots value the function is unbounded?

OpenStudy (kinggeorge):

Since \[|\sin(x)| \leq 1\]You know that \[|3\;\sin^4(2x)|\leq 3\]Thus, \[1+3\;\sin^4(2x) \leq4\]Likewise, you also know that \[1+3\;\sin^4(2x) \geq -2\]Since the range of this function is bounded within the interval \([-2, 4]\), the function is bounded.

OpenStudy (anonymous):

how did you find the range so quickly?

OpenStudy (anonymous):

yes, where did you get -2?

OpenStudy (anonymous):

did you use this tirg identity sin^2(2x) - 1/2(1-cos(2x)) ?

OpenStudy (anonymous):

sin^2(2x) = 1/2(1-cos(2x))

OpenStudy (kinggeorge):

I got the -2 from \[|3\;\sin^4(2x)|\leq3\Rightarrow -3\leq 3\;\sin^4(2x)\Rightarrow -2\leq1+3\;\sin^4(2x)\]However, I realize now that you can restrict it to \(+1\) instead of \(-2\).

OpenStudy (anonymous):

i understand that 3 is the amplitude, and then you add one to get minimum of -2, but wouldnt the exponent have an effect?

OpenStudy (anonymous):

yeah, i got 1

OpenStudy (kinggeorge):

@mridrik you are correct. Because you have a \(\sin^4(x)\), it must always be positive, since any negative number raised to an even power will be positive.

OpenStudy (anonymous):

thanks @ all

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

this website is really awesome because I am self teaching myself & whenever I am stuck and google doesn't help then someone later or sooner resolves my problem

OpenStudy (anonymous):

yeah, i agree, kindgeorge, what would the range be if sin was taken to the 3rd power instead?

OpenStudy (anonymous):

kinggeorge*

OpenStudy (kinggeorge):

If it was to the third, it would be at -2 instead of +1. If the 3 was contained in the power it would be -26.

OpenStudy (anonymous):

? -26? where'd you get that?

OpenStudy (kinggeorge):

\((-3)^3=-27\) so if you add one you get \(-26\)

OpenStudy (anonymous):

so the minimum is -26?

OpenStudy (anonymous):

what is the proceadure for finding the range? (especially in this case)

OpenStudy (kinggeorge):

If the function were instead \[1+(3\sin(2x))^3\]the minimum would be \(-26\)

OpenStudy (anonymous):

i meant 3sin^3, but continue

OpenStudy (kinggeorge):

To find the range (in this case), first find the range of \(\sin(x)\) This is \([-1, 1]\)Then, if you have an even exponent, the range become \([0, 1]\). If you have an odd exponent, nothing happens. Once you have that, multiply the range by any number in front of the sin. In this case you have a 3, so the range goes to \([0*3, 1*3]=[0, 3]\). Finally, add/subtract any constant in the front so in this case it becomes \([1, 4]\).

OpenStudy (anonymous):

Well explained!

OpenStudy (kinggeorge):

@mridrik the range of \[1+3\;\sin^3(2x)\]would be \([-2, 4]\)

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

what is the method to find the domain?

OpenStudy (kinggeorge):

To find the domain, just find the numbers the function is defined on. For this function, the domain would be \((-\infty, \infty)\) because all the real numbers are defined for \(2x\), \(\sin(x)\), \(x^4\), \(3x\) and \(1+x\) (Your function uses all of these functions). If one of these wasn't defined for all numbers, then the domain of your function would not include those numbers it's being divided by. If, for example, your function was \[1+{3\;\sin^4 (2x)\over x}\]Then the domain would be \((-\infty, 0), (0, \infty)\) because \(x=0\) isn't defined for \(1\over x\).

OpenStudy (anonymous):

thankssssssssssss

OpenStudy (kinggeorge):

You're welcome.